OpenStudy (anonymous):

evaluate 8x-4y when x=5 and y=2

OpenStudy (britanyashton):

substitute x and y into the problem.

OpenStudy (britanyashton):

then solve.

OpenStudy (yamyam70):

@trinajoy should be good for you :)

OpenStudy (anonymous):

ok so how u are going to work this out is

OpenStudy (anonymous):

what are X and Y

OpenStudy (anonymous):

ok.. X and Y are 5 and 2 right

OpenStudy (anonymous):

solve 85-42 u replace the X and Y with the numbers that tell u (X=5 & Y= 2)

OpenStudy (anonymous):

if u need questions please ask

OpenStudy (anonymous):

I mean if u have any questions please ask sorry

OpenStudy (korosh23):

Are you trying to find the answer to the equation above?

OpenStudy (anonymous):

8x5=45 and 4x2=8 so it's 45+8=53. :) hope I helped!!

OpenStudy (anonymous):

To solve this 8(x=5)-2(y=2) so the problem looks like 8(5)-4(2) = 40-8 =32 The answer is 32

OpenStudy (anonymous):

Okay people are answering this wrong so I wanted to post it is 8 times five minus 4 times two which would be 40-8 which would be 32. Please vote my answer as best so I get extra credit in my math class (: Thank you!